Calculus II, Spring 2019 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 07

1. Section 14.2: Problems 41, 49, 61, 63. Hint for problems 41, 61, 63: Read ‘Changing to Polar Coordinates’ on page 781. Note that in equation (1), page 781, “|r| < δ” should be “0 < |r| < δ” instead. Hint for problem 49: Find two paths that pass through (1, 1). Remark: No recitation, no quiz next week. This homework will be covered in the recitation on Monday April 15.